OpenStudy (anonymous):

Simplify. √108-3√8-√27

jhonyy9 (jhonyy9):

108=4*27=4*9*3 so than sqrt108 =6sqrt3 3sqrt8 =6sqrt2 sqrt27 =3sqrt3 sqrt108 -3sqrt8 -sqrt27 =6sqrt3 -6sqrt2 -3sqrt3 = =3sqrt3 -6sqrt2 =3(sqrt3 -2sqrt2)

OpenStudy (anonymous):